In a triangle ABC, AD is median to side BC. Angle BDA=45 degree and CAD = 30 degree. Then find angle BAD.

Grade:9

1 Answers

Sarah Saha
8 years ago
Given angle BDA=45 degree &angle CAD=30 degree
      Now ,In tri.ADC,angle ADC =(180-45)=135 degree & angles (ADC+CAD+ACD)=180 degree
       From this,we get angle ACD=15 degree.
  Now,BD=CD,i.e BC=2CD i.e angle BAC=2CAD i.e angle BAC=2x30=60
 Therefore;angle BAD=(BAC-CAD)=60-30=30 degree